# Conditional Prototype Rectification Prompt Learning

### v1

@@ 4.3 Tuning for Downstream Tasks
Table 1: Comparison with existing methods in the base-to-new generalization setting with ViT-B/16 as the backbone. The context length M is 4 for prompot-based methods with the 16-shots samples from the base classes. H: Harmonic mean.
(a) Average over 11 datasets
Base New H
CLIP 69.34 74.22 71.70
CoOp 82.63 67.99 74.60
CoCoOp 80.47 71.69 75.83
ProGrad 82.48 70.75 76.16
KgCoOp 80.73 73.60 77.00
MaPLe 82.28 75.14 78.55
CPR 83.81 76.21 79.82
(b) ImageNet
Base New H
CLIP 72.43 68.14 70.22
CoOp 76.46 66.31 71.02
CoCoOp 75.98 70.43 73.10
ProGrad 77.02 66.66 71.46
KgCoOp 75.83 69.96 72.78
MaPLe 76.66 70.54 73.47
CPR 77.21 70.60 73.75
(c) Caltech101
Base New H
CLIP 96.84 94.00 95.40
CoOp 98.11 93.52 95.76
CoCoOp 97.96 93.81 95.84
ProGrad 98.02 93.89 95.91
KgCoOp 97.72 94.39 96.03
MaPLe 97.74 94.36 96.02
CPR 98.64 95.41 97.00
(d) OxfordPets
Base New H
CLIP 91.17 97.26 94.12
CoOp 94.24 96.66 95.43
CoCoOp 95.20 97.69 96.43
ProGrad 95.07 97.63 96.33
KgCoOp 94.65 97.76 96.18
MaPLe 95.43 97.76 96.58
CPR 95.48 97.77 96.61
(e) StanfordCars
Base New H
CLIP 63.37 74.89 68.65
CoOp 76.20 69.14 72.49
CoCoOp 70.49 73.59 72.01
ProGrad 77.68 68.63 72.88
KgCoOp 71.76 75.04 73.36
MaPLe 72.94 74.00 73.47
CPR 77.39 75.12 76.24
(f) Flowers102
Base New H
CLIP 72.08 77.80 74.83
CoOp 97.63 69.55 81.23
CoCoOp 94.87 71.75 81.71
ProGrad 95.54 71.87 82.03
KgCoOp 95.00 74.73 83.65
MaPLe 95.92 72.46 82.56
CPR 96.96 79.08 87.11
(g) Food101
Base New H
CLIP 90.10 91.22 90.66
CoOp 89.44 87.50 88.46
CoCoOp 90.70 91.29 90.99
ProGrad 90.37 89.59 89.98
KgCoOp 90.50 91.70 91.09
MaPLe 90.71 92.05 91.38
CPR 90.95 92.03 91.48
(h) FGVCAircraft
Base New H
CLIP 27.19 36.29 31.09
CoOp 39.24 30.49 34.30
CoCoOp 33.41 23.71 27.74
ProGrad 40.54 27.57 32.82
KgCoOp 36.21 33.55 34.83
MaPLe 37.44 35.61 36.50
CPR 41.00 35.69 38.16
(i) SUN397
Base New H
CLIP 69.36 75.35 72.23
CoOp 80.85 68.34 74.07
CoCoOp 79.74 76.86 78.27
ProGrad 81.26 74.17 77.55
KgCoOp 80.29 76.53 78.36
MaPLe 80.82 78.70 79.75
CPR 82.38 79.44 80.89
(j) DTD
Base New H
CLIP 53.24 59.90 56.37
CoOp 80.17 47.54 59.68
CoCoOp 77.01 56.00 64.85
ProGrad 77.35 52.35 62.45
KgCoOp 77.55 54.99 64.35
MaPLe 80.36 59.18 68.16
CPR 82.87 58.45 68.55
(k) EuroSAT
Base New H
CLIP 56.48 64.05 60.03
CoOp 91.54 54.44 68.27
CoCoOp 87.49 60.04 71.21
ProGrad 90.11 60.89 72.67
KgCoOp 85.64 64.34 73.48
MaPLe 94.07 73.23 82.35
CPR 92.71 74.18 82.41
(l) UCF101
Base New H
CLIP 70.53 77.50 73.85
CoOp 85.14 64.47 73.37
CoCoOp 82.33 73.45 77.64
ProGrad 84.33 74.94 79.35
KgCoOp 82.89 76.67 79.65
MaPLe 83.00 78.66 80.77
CPR 86.35 79.83 82.96
5 Experiment
5.1 Datasets and Implementation Details
Similar to CoCoOp [29] and ProGrad [31], our evaluation of the proposed method encompasses two primary settings: 1) generalization from base to new classes within a dataset; and 2) few-shot image classification. These evaluations are conducted leveraging the pretrained CLIP [19] model, ensuring a consistent basis for comparison and thorough assessment of generalization and adaptability capabilities. More detailed results will be given in the Supplementary materials.
Datasets. Following CLIP [19], CoOp [30], CoCoOp [29], and ProGrad [31], the base-to-new generaliation and few-shot image classification are conducted on 11 image classification datasets, i.e., ImageNet [6] and Caltech101 [8] for generic object classification; OxfordPets [18], StanfordCars [14], Flowers102 [17], Food101 [2], and FGVCAircraft [16] for fine-grained visual categorization, EuroSAT [11] for satellite image classification, UCF101 [22] for action recognization, DTD [5] for texture classification, and SUN397 [25] for scene recognition.
Training details. Our experiments are executed using the same frameworks as established in CoOp [30] and ProGrad [31], incorporating the CLIP model as the core component. The experiments utilize ResNet-50 [10] and ViT-B/16 [7] as vision backbones, aligning with the setups in CoOp where the context length is consistently held at 16. To ensure statistical reliability and fairness in comparison, the reported performance metrics are the averages derived from three distinct random seeds. Consistency with the original studies also extends to the training scheme; we adhere to the same training epochs, training schedule, and data augmentation strategies as described in CoOp [30] and ProGrad [31]. Regarding hyperparameters, we set the $\alpha$ for the base-to-new generalization and the few-shot learning setups to 8.0 and 1.0, respectively, and uniformly set alpha ($\alpha$) to 0.95, and we adjust the number of neighbors ($k$) to 5. The computational groundwork for all experiments is provided by NVIDIA RTX A100 80G GPUs, ensuring robust and efficient processing capabilities for our analyses.
Baselines. Four type of CoOp-based methods and three adapter-style methods are treated as baselines for comparison: CoOp [30], CoCoOp [29], ProGrad [31], KgCoOp [26], MaPle [13], CLIP-Adapter [9], TaskRes [27] and GraphAdapter [15].
@@ 5.2 Few-shot Classification
5.3 Generalization From Base-to-New Classes
Similar to previous works [30, 29, 31, 26, 13], we divide each dataset into two categories: base classes (Base) and new classes (New). Mirroring the zero-shot setting, the new classes are disjoint from the base classes. To assess the generalizability of the CoOp-based approaches, all evaluated methods, including the proposed CPR, utilize the base classes for tuning and perform evaluations on the new classes. The specific results are detailed in Table 1, which provides comprehensive performance metrics across all 11 datasets using the ViT-B/16 backbone with 16-shot samples.
As shwon in Table 1, CPR method surpasses existing methodologies in achieving a higher average performance based on the Harmonic mean across all settings, underlining its enhanced capability for generalizing from base to new classes. Specifically, CPR demonstrates superior performance on new classes in nine out of the eleven datasets, including ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, FGVCAircraft, SUN397, EuroSAT, and UCF101. The exceptional performance highlights the improvement of CPR in generalization to new classes compared to the current methods. Moreover, CPR consistently achieves top results for the base classes in the majority of scenarios. Hence, CPR effectively enhances generalizability to new classes while maintaining or even improving base class performance, thereby securing the highest Harmonic mean across all evaluated datasets.
5.4 Ablation Studies
The effects of each prototype rectification strategy in CPR. To evaluate the effectiveness of the proposed strategy, ablation studies were conducted. As shown in Table 2, the results indicate that both strategies significantly enhanced the generalizability of model and mitigated overfitting relative to the base class. If the CoAdapter is removed from CPR, the accuracy on new class recognition tasks would decrease by 5.5%. Similarly, if NNR is not used, the accuracy would decrease by 5.9%. These research findings confirm that each strategy can independently enhance recognition performance, and their synergistic application is particularly effective in improving accuracy in efficient transfer learning scenarios.
@@ 5.4 Ablation Studies
The effects of neighbors in NNR. In the Nearest Neighbor Rectification (NNR) strategy, determining the optimal value of $k$—the number of nearest unlabeled samples to associate with each fused prototype $P_{i}$-is essential for knowledge completion across each class. To ascertain the most appropriate $k$ value, we conducted an ablation study within a few-shot learning framework, evaluating $k$ values from the set ${1,3,5,7,9}$. The results, depicted in Figure 4, reveal that the model attains its best performance when $k$ is set to 5. This outcome implies that, within the context of our experiments, integrating information from five nearest neighbors achieves the optimal balance between acquiring sufficient knowledge and mitigating the influence of irrelevant information, thereby enhancing the predictive accuracy of model most effectively.
@@
The effects of $\alpha$ in NNR. In the Nearest Neighbor Rebalancing (NNR) strategy, the parameter $\alpha$ plays a pivotal role in determining the extent to which the model benefits from the knowledge in unlabeled data. To identify the ideal setting for $\alpha$, we carried out a series of ablation experiments. As shown in Figure 4, the results indicate that the model achieves optimal performance when $\alpha$ is set to 0.95. This outcome suggests that an appropriate $\alpha$ value can effectively balance the insights from unlabeled data with the requirements to ensure the reliability of the correction process. In contrast, too high of an $\alpha$ value might not fully utilize the embedded knowledge in the unlabeled data, while too low of an $\alpha$ might introduce unnecessary noise or irrelevant details, potentially impairing the performance.
@@
The effects of different backbones. As shown in Table 4, we also evaluate the effectiveness of our Conditional Prototype Rectification Prompt Learning (CPR) method across various CLIP visual backbones, namely ResNet-50, ResNet-101, ViT-B/32, and ViT-B/16. Across all four visual backbones, CPR consistently outperforms prior methods, demonstrating its robustness and adaptability to different underlying architectures. This result highlights versatility of CPR and its ability to leverage distinct visual features extracted by various backbones, thus affirming its broad applicability and effectiveness in enhancing model performance.
@@
The effects of $\lambda$. To investigate the effectiveness of hyperparameter $\lambda$, as specified in Equation (8) in Section 4.3, we conduct experiments on FGVCAircraft by varying different $\lambda$. $\mathcal{L}_{cons}$ aims to enhance the generalizability for unseen classes by minimizing the distance between learnable fused features and fixed textual enhanced representations provided by the LLM As shown in Table 4, selecting an optimal value for $\lambda$ is crucial for enhancing model performance. Specifically, for base-to-new generalization, setting $\lambda$ to 8 yields the most beneficial outcome. An appropriate $\lambda$ can avoid learning text tokens with poor generalization on base classes while generating sample-specific representations. On the other hand, for few-shot learning setting, the optimal performance is observed when $\lambda$ is adjusted to 1.
@@
Limitations. The limitation of our CPR is rooted in the selection process of unlabeled data. In our current methodology, we employ a $k$-nearest neighbor algorithm to identify supplementary knowledge. While this strategy provides a foundational approach to augmenting our dataset, it lacks a definitive level of confidence, potentially limiting the quality of the knowledge integration. We posit that a more precise and sophisticated selection method could significantly enhance the quality of the supplementary knowledge, thereby improving the performance on downstream tasks.
@@
6 Conclusion
In this paper, we provide an in-depth examination of the constraints faced by previous approaches in scenarios characterized by scarce data, identifying two main limitations: 1) the reliance on a single modality for modeling task-specific knowledge, and 2) the employment of resource-intensive and protracted techniques for knowledge enhancement, which are crucial for tasks requiring data efficiency. To address these issues, we introduce an innovative Conditional Prototype Rectification Prompt Learning (CPR) methodology for visual-language models. CPR aims to rectify biases present in base examples and effectively enrich limited datasets. It achieves this by amalgamating fused textual and visual prototype knowledge with insights gleaned from unlabeled samples, thereby enabling the classifier to adapt more effectively to downstream tasks. Our comprehensive testing across 11 benchmark datasets demonstrates the effectiveness of our CPR on few-shot learning and generalization.

### v2

@@ Introduction
Table 1: Comparison with existing methods in the base-to-new generalization setting with ViT-B/16 as the backbone. The context length M is 4 for prompot-based methods with the 16-shots samples from the base classes.
Experiment
Datasets and Implementation Details
Similar to CoCoOp (Zhou et al. 2022a) and PromptSRC (Khattak et al. 2023b), our evaluation of the proposed method encompasses two primary settings: 1) generalization from base to new classes within a dataset; and 2) few-shot image classification. These evaluations are conducted leveraging the pretrained CLIP (Radford et al. 2021) model, ensuring a consistent basis for comparison and thorough assessment of generalization and adaptability capabilities. More detailed results will be given in the Supplementary materials.
Datasets. Following CoOp (Zhou et al. 2022b), CoCoOp (Zhou et al. 2022a), and TCP (Yao, Zhang, and Xu 2024), the base-to-new generaliation and few-shot image classification are conducted on 11 image classification datasets, i.e., ImageNet (Deng et al. 2009) and Caltech101 (Fei-Fei, Fergus, and Perona 2007) for generic object classification; OxfordPets (Parkhi et al. 2012), StanfordCars (Krause et al. 2013), Flowers102 (Nilsback and Zisserman 2008), Food101 (Bossard, Guillaumin, and Van Gool 2014), and FGVCAircraft (Maji et al. 2013) for fine-grained visual categorization, EuroSAT (Helber et al. 2019) for satellite image classification, UCF101 (Soomro, Zamir, and Shah 2012) for action recognization, DTD (Cimpoi et al. 2014) for texture classification, and SUN397 (Xiao et al. 2010) for scene recognition.
Training details. Our experiments are executed using the same frameworks as established in TCP (Yao, Zhang, and Xu 2024) and GraphAdapter (Li et al. 2023), incorporating the CLIP model as the core component. The experiments utilize ResNet-50 (He et al. 2016) and ViT-B/16 (Dosovitskiy et al. 2021) as vision backbones, aligning with the setups in CoOp where the context length is consistently held at 16. To ensure statistical reliability and fairness in comparison, the reported performance metrics are the averages derived from three distinct random seeds. Consistency with the original studies also extends to the training scheme; we adhere to the same training epochs, training schedule, and data augmentation strategies as described in CoOp (Zhou et al. 2022b). Regarding hyperparameters, we set the $\alpha$ for the base-to-new generalization and the few-shot learning setups to 8.0 and 1.0, respectively, and uniformly set $\alpha$ to 0.95, and we adjust the number of neighbors $k$ to 5. The computational groundwork for all experiments is provided by a single NVIDIA RTX A100 80G GPU.
@@
Generalization From Base-to-New Classes
Similar to previous works (Yao, Zhang, and Xu 2024; Khattak et al. 2023b, a), we divide each dataset into two categories: base classes and new classes. Mirroring the zero-shot setting, the new classes are disjoint from the base classes. To assess the generalizability of the CoOp-based approaches, all evaluated methods, including the proposed CPR, utilize the base classes for tuning and perform evaluations on the new classes. The specific results are detailed in Table 1, which provides comprehensive performance metrics across all 11 datasets using the ViT-B/16 backbone with 16-shot samples.
As shwon in Table 1, CPR method surpasses existing methodologies in achieving a higher average performance based on the harmonic mean across all settings, underlining its enhanced capability for generalizing from base to new classes. Specifically, CPR demonstrates superior performance on new classes in seven out of the eleven datasets, including ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, Food101 and SUN397. The exceptional performance highlights the improvement of CPR in generalization to new classes compared to the current methods. Moreover, CPR consistently achieves top results for the base classes in the majority of scenarios. Hence, CPR effectively enhances generalizability to new classes while maintaining or even improving base class performance, thereby securing the highest Harmonic mean across all evaluated datasets.
Ablation Studies
The effects of each prototype rectification strategy in CPR. To evaluate the effectiveness of the proposed strategy, ablation studies were conducted. As shown in Table 2, the results indicate that both strategies significantly enhanced the generalizability of model and mitigated overfitting relative to the base class. If the CoAdapter is removed from CPR, the accuracy on new class recognition tasks would decrease by 5.5%. Similarly, if NNR is not used, the accuracy would decrease by 5.7%. These research findings confirm that each strategy can independently enhance recognition performance, and their synergistic application is particularly effective in improving accuracy in efficient transfer learning scenarios.
@@
The effects of neighbors in NNR. In the NNR strategy, determining the optimal value of $k$—the number of nearest unlabeled samples to associate with each fused prototype $P_{i}$-is essential for knowledge completion across each class. To ascertain the most appropriate $k$ value, we conducted an ablation study within a few-shot learning framework, evaluating $k$ values from the set $\{1,3,5,7,9\}$. The results, depicted in Figure 4, reveal that the model attains its best performance when $k$ is set to 5. This outcome implies that, within the context of our experiments, integrating information from five nearest neighbors achieves the optimal balance between acquiring sufficient knowledge and mitigating the influence of irrelevant information, thereby enhancing the model’s predictive accuracy most effectively.
@@
The effects of $\alpha$ in NNR. In the NNR strategy, the parameter $\alpha$ plays a pivotal role in determining the extent to which the model benefits from the knowledge in unlabeled data. To identify the ideal setting for $\alpha$, we carried out a series of ablation experiments. As shown in Figure 4, the results indicate that the model achieves optimal performance when $\alpha$ is set to 0.95. This outcome suggests that an appropriate $\alpha$ value can effectively balance the insights from unlabeled data with the requirements to ensure the reliability of the correction process. In contrast, too high of an $\alpha$ value might not fully utilize the embedded knowledge in the unlabeled data, while too low of an $\alpha$ might introduce unnecessary noise or irrelevant details, potentially impairing the performance.
@@
The effects of different backbones. As shown in Table 4, we also evaluate the effectiveness of our CPR method across various CLIP visual backbones, namely ResNet-50, ResNet-101, ViT-B/32, and ViT-B/16. Across all four visual backbones, CPR consistently outperforms prior methods, demonstrating its robustness and adaptability to different underlying architectures. This result highlights CPR’s versatility and its ability to leverage distinct visual features extracted by various backbones, thus affirming its broad applicability and effectiveness in enhancing model performance.
@@
The effects of $\lambda$. To investigate the effectiveness of hyperparameter $\lambda$, as specified in Equation (8) in Section Approch, we conduct experiments on FGVCAircraft by varying different $\lambda$. $\mathcal{L}_{cons}$ aims to enhance the generalizability for unseen classes by minimizing the distance between learnable fused features and fixed textual enhanced representations provided by the LLM As shown in Table 5, selecting an optimal value for $\lambda$ is crucial for enhancing model performance. Specifically, for base-to-new generalization, setting $\lambda$ to 8 yields the most beneficial outcome. An appropriate $\lambda$ can avoid learning text tokens with poor generalization on base classes while generating sample-specific representations. On the other hand, for few-shot learning setting, the optimal performance is observed when $\lambda$ is adjusted to 1.
@@
Limitations. The limitation of our CPR is rooted in the selection process of unlabeled data. While the $k$-nearest neighbor algorithm provides a foundational approach to augmenting our dataset, it lacks a definitive level of confidence, potentially limiting the quality of the knowledge integration. We posit that a more precise and sophisticated selection method could significantly enhance the quality of the supplementary knowledge, thereby improving the performance on downstream tasks.
@@
Conclusion
In this paper, we provide an in-depth examination of the constraints faced by previous approaches in scenarios characterized by scarce data, identifying two main limitations: 1) the reliance on a single modality for modeling task-specific knowledge, and 2) the employment of resource-intensive and protracted techniques for knowledge enhancement, which are crucial for tasks requiring data efficiency. To address these issues, we introduce an innovative Conditional Prototype Rectification Prompt Learning (CPR) methodology for visual-language models. CPR aims to rectify biases present in base examples and effectively enrich limited datasets. It achieves this by amalgamating fused textual and visual prototype knowledge with insights gleaned from unlabeled samples, thereby enabling the classifier to adapt more effectively to downstream tasks. Our comprehensive testing across 11 benchmark datasets demonstrates the effectiveness of our CPR on few-shot learning and generalization tasks.